q/x+p/4x=2. Solve for x.

Answer by Alan3354(69443)

q/x+p/4x=2. Solve for x.
--------
(4q/4x) + (p/4x) = 2
(4q + p)/(4x) = 2
8x = 4q + p
x = (4q + p)/8
-------------------
Was that hard?